Use the Division Algorithm to find the greatest common divisor of each pair of numbers below and determine whether each pair is rela- tively prime or not. Then reverse the process and write the gcd as a sum of multiples of the original pair.

a. 12 and 15 b. 36 and 72 c. 27 and 10 d. 35 and 12
Answer #1


@ By division algorithm 15- 12x1 + 3 12= 3x4 to Iged (15.12) = 3 Now, 3 = 1615) + F1112) ③ By division algorithm - 72 = 36*2